GachaGirl23:

The scale model of a rectangular patio will be enlarged by a scale factor of 2.5 to create the actual patio. The actual patio will be 15 feet by 20 feet. Which is the area of the scale model?

GachaGirl23:

48 square feet 60 square feet 120 square feet 300 square feet

GachaGirl23:

option choices

Convert:

Well I think for this one you would divide each number by 2.5 to find what the scale model was: 15/2.5=6 20/2.5=8 So based on this we know that the scale model was 6 by 8 As we know, area is length times width. 6*8=area 6*8=48 Area=48 feet So 48 feet would be your answer.

Convert:

So because A is 48 square feet, that is what you would choose.

GachaGirl23:

okay thank you for figuring it out and even bothering to check out my question!

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @GachaGirl23 okay thank you for figuring it out and even bothering to check out my question! \(\color{#0cbb34}{\text{End of Quote}}\) I'm happy to help! :)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert Well I think for this one you would divide each number by 2.5 to find what the scale model was: 15/2.5=6 20/2.5=8 So based on this we know that the scale model was 6 by 8 As we know, area is length times width. 6*8=area 6*8=48 Area=48 feet So 48 feet would be your answer. \(\color{#0cbb34}{\text{End of Quote}}\) good job !

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert Well I think for this one you would divide each number by 2.5 to find what the scale model was: 15/2.5=6 20/2.5=8 So based on this we know that the scale model was 6 by 8 As we know, area is length times width. 6*8=area 6*8=48 Area=48 feet So 48 feet would be your answer. \(\color{#0cbb34}{\text{End of Quote}}\) good job ! \(\color{#0cbb34}{\text{End of Quote}}\) Thank you jhonyy (: